Solve the system by substitution.
x + y = -7
x - y = 13

Found 2 solutions by adamchapman, fractalier:
Answer by adamchapman(301)

x+y=-7 ..1
x-y=13 ..2
rearrange equation(..2) to give:
y=x-13 ..3
now substitute equation (..3) into equation (..1) for y:
x+(x-13)=-7
2x-13=-7
2x=6
x=3 ..4
now substitute equation (..4) into equation (..3) for y:
y=3-13
y=-10
Therefore x=3 and y=-10
I hope this helps. If you have any other problems, please dont hesitate to contact me at adam.chapman@student.manchester.ac.uk

Answer by fractalier(6550)

From
x + y = -7
x - y = 13
let us solve the first one for x and substitute it into the second...thus we have
x = -y - 7 and
(-y - 7) - y = 13
-y - 7 - y = 13
-2y = 20
y = -10 and then
x = 3